OpenStudy (anonymous):

what is the integral of (4-x-y)dx.

OpenStudy (slaaibak):

\[4x - {x^2 \over 2} - yx + C\]

OpenStudy (zarkon):

I would write \[4x-\frac{x^2}{2}-yx+C(y)\] where \(C(y)\) is just a function of \(y\) since \(\displaystyle \frac{\partial}{\partial x}\left[4x-\frac{x^2}{2}-yx+C(y)\right]=4-x-y\)